# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v4

@@ 2 Related Works
0-shot Linear
CLIP L/14-336 76.6 84.8
+LLM2CLIP 74.1 85.2
CLIP L/14 75.5 83.9
+LLM2CLIP 71.8 84.4
CLIP B16 70.2 79.9
+LLM2CLIP 67.3 80.6
Table 4: Imagenet 0-shot classification and linear probe results.
MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
@@
Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold. We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
Methods Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
Qwen2.5-0.5B-CC 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
Llama-3.2-1B-CC 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Llama-3-8B-CC 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
DeepSeek-R1-Distill-Llama-8B-CC 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
Llama-3.1-8B-CC 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
@@
Table 6: Ablation for using different text encoder. BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs. The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
@@
Table 7: Ablation study on the proportion of dense captions from different MLLMs. Here, we explore the impact of both genuine image captions and fake dense captions generated by MLLMs on the performance of  LLM2CLIP.
Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@
89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
Linear(×$\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
Linear(×$\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
Linear(×$\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Linear(×$\times 4$) Linear(×$\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
Transformer(×$\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
Transformer(×$\times 1$) Transformer(×$\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
[…]
3 Methods
3.1 Stage 1: LLM Caption Contrastive Fine-tuning
Traditionally, LLM output layers serve as classification heads producing text tokens. Prior works, such as LLM2Vec [2] and NV-EMBED-v2 [17], have explored converting LLMs into universal text encoders by applying additional designs and fine-tuning. Inspired by these approaches, we adapt similar concepts to the novel scenario of cross-modal pretraining. Our key design considerations encompass model architecture, training methods, and training data, summarized as follows:
Model Architecture:
  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability. We experimented with two adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, and (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs. However, neither approach significantly boosted CLIP performance during embedding fine-tuning. Consequently, we opted not to use adaptors during the LLM caption contrastive fine-tuning phase for simplicity.
Training Methods:
  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens. Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention. While effective in pure text tasks (as shown by LLM2Vec), MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
Training Data: We conducted experiments with Dreamlip [41]’s caption dataset, naturally containing multiple captions per image, making it ideal for supervised textual contrastive learning. Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings [29] into our contrastive training process. 89213
3.2 Stage 2: LLM2CLIP Post Fine-tuning
We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining. By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective post-training step. CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model. Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and determining the optimal approach to enhance CLIP performance is critical. Here, we elaborate on model architecture, training methods, and training data.
Model Architecture:
  • Fine-tuning Parameters and Adaptor Design: Training CLIP typically requires large-scale datasets and significant batch sizes, resulting in considerable computational costs. Therefore, we aim to employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this post-training stage. By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge. Given the large size of the LLM, we compare two fine-tuning strategies: (1) directly fine-tuning the LLM using LoRA, and (2) freezing the LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain the sentence token features. Method (2) offers distinct advantages, including completely eliminating gradient updates for the LLM, thus significantly reducing GPU memory consumption. Moreover, since the adaptor is placed after the LLM, we can even precompute and store text features through offline inference, thereby reducing the inference overhead from multiple epochs to just a single pass. Consequently, this approach eliminates the need to load the LLM into GPU memory during training, greatly accelerating experiment iteration and training time, and drastically reducing GPU memory footprint. By default, we adopt method (2) with an adaptor consisting of four linear blocks. Detailed efficiency analysis will be presented in Section 4.4.
Training Methods: We tested several approaches:
  • (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
  • (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
  • (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
  • (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
Extensive experiments (see Table 10) led us to adopt method (a) for its simplicity and effectiveness.
Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions. Datasets such as DreamLIP and Recaption, which leverage Share-Captioner [3], InstructBLIP [5] and LLaVA-1.5 [23] to generate multiple dense captions per image, can be effectively utilized. We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions. Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
Stage 1 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
@@ 3.2 Stage 2: LLM2CLIP Post Fine-tuning
Table 9: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.

### v5

@@ 4.2 System Comparison
0-shot* 0-shot Linear
CLIP L/14-336 74.9 76.6 84.8
+LLM2CLIP 74.6 75.8 85.2
CLIP L/14 73.7 75.5 83.9
+LLM2CLIP 73.2 74.3 84.4
Table 3: Zero-shot classification and linear probe performance on ImageNet. *0-shot uses the class template ‘a photo of the {classname}’ only.
Method Zero-shot Seg. mIOU OV-COCO Det. COCO val2017
COCO-S ADE VOC City Novel Base All APbb/APseg
EVA02 12.9 11.5 21.0 13.5 24.7 53.6 46.0 45.0/38.2
+LLM2CLIP 15.3 15.8 29.1 20.1 28.9 54.7 48.0 45.6/38.7
Table 4: Zero-shot/supervised segmentation, open-vocabulary detection benchmarks. COCO val2017 is supervised results using CLIP’s visual encoder.
Model VQA Pope $M$ Seed
V2 GQA Vz SQA TV R A P MME MB MC LB MV All I V
@@
Table 5: Performance of Llava 1.5 benchmarks. For + LLM2CLIP we replace Llava’s CLIP Vit-L/14 with our finetuned version.
Fine-tuning Data Volume Analysis. In Table 1, we experimented with 3M, 15M, and 60M training data on CLIP and EVA02 using ViT-L/14. More data consistently improves retrieval performance for both long and short text tasks at 224 and 336 resolutions. With 3M data, EVA02 shows noticeable improvement in long text retrieval but minimal gain in short text retrieval, suggesting the model first compensates for disrupting the original cross-modal space before achieving broader improvements.
Zero-Shot Multilingual Text-Image Retrieval As in Table 2, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLIP2.
Zero-shot & Linear Probe ImageNet Classification We evaluate our model on ImageNet using two standard protocols: (1) Zero-shot classification, where we report accuracy using the prompt template “a photo of the {classname}”, following the CLIP methodology. We also evaluate the average performance over 80 handcrafted prompt variants, as commonly practiced in prior work. (2) Linear probing, where we freeze the visual encoder and train a linear classifier using fixed hyperparameters across all experiments (batch size 1024, learning rate 0.1, momentum 0.9, weight decay 0, 50 epochs with SGD).
As shown in Table 3, we conduct experiments using CLIP ViT-L/14 at both 224 and 336 resolutions. We observe that after applying LLM2CLIP, zero-shot performance on ImageNet shows a modest drop. However, the accuracy from linear probing improves over the original CLIP baseline. This suggests that while the quality of the learned visual features remains strong (as linear probing only tests CLIP visual encoder’s representation power using a supervised head), the overall multimodal alignment in the shared space may slightly deteriorate since zero-shot classification drops, especially for fine-grained category separation.
We hypothesize that the drop in zero-shot performance could be mitigated by increasing the amount of LLM2CLIP fine-tuning data, which remains to be verified in future work. Nonetheless, this trade-off echoes a common pattern observed in recent CLIP fine-tuning studies, such as LongCLIP (Zhang et al. 2024a) and CLIP-MoE (Zhang et al. 2024b), where performance on head classes improves while long-tail categories suffer due to distribution imbalance—particularly under limited-scale fine-tuning. The contrast with retrieval results is instructive: retrieval benefits greatly because its vocabulary is broad and frequent, whereas classification demands uniform discriminability across all (often obscure) class nouns.
Zero-Shot / Supervised Segmentation and Object Detection As our linear-probe results already suggest, the visual features produced by the encoder become noticeably stronger after LLM2CLIP training. Table 4 further confirms this on both zero-shot and fully supervised object detection and segmentation benchmarks. In the zero-shot setting—which relies on the text encoder—LLM2CLIP improves CLIP’s multimodal representations; in the supervised setting—where only the vision encoder is finetuned—it also enhances the purely visual, low-level feature extractor. We conjecture that these gains stem from the LLM’s ability to parse dense captions more accurately, especially spatial terms and object-to-object relations, and to inject that knowledge into the joint vision–language space.
Multimodal Large Language Models Performance Following LLAVA1.5 (Liu et al. 2023b), we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction. As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving CLIP visual encoder’s abilities for complex image reasoning and understanding.
Stage1 Avg
I2T T2I
Lora, AvgPool, Bidirectional, Supervise Simcse 80.4 77.9
Lora, AvgPool, Bidirectional, Un-supervise Simcse 59.2 57.7
Lora, AvgPool, Bidirectional, MNTP 70.1 67.0
Lora, AvgPool, Bidirectional, MNTP, Supervise Simcse 79.7 77.2
Lora, AvgPool, Casual, Supervise Simcse 80.0 77.5
Lora, EOS, Bidirectional, Supervise Simcse 80.0 77.3
Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 74.1 71.3
Table 6: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. red text indicates content that differs from the default setting.
Method Avg
I2T T2I
CLIP 74.4 72.0
Directly Finetune (50%) 74.5 72.3
bge-en-icl 78.9 78.2
LLM2Vec-Llama-3-8B 81.4 80.2
NV-Embed-v2 81.4 79.9
VLM2Vec 78.2 77.1
bge-m3-XLM-R 65.0 63.6
jina-v3-XLM-R 73.6 71.0
e5 (XLM-R) 74.0 71.7
Qwen2.5-0.5B-$C$ 75.6 73.0
Llama-3.2-1B-$C$ 80.4 77.9
Llama-3-8B-$C$ 83.4 80.9
DeepSeek-R1-Distill-Llama-8B-$C$ 83.5 80.5
Llama-3.1-8B-$C$ 84.8 81.0
Llama3.1-8B 66.5 62.5
@@
Table 7: Ablation for using different text encoder. ”-CC” indicates encoders with caption contrastive fine-tuning.
4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section. We use a subset of the CC 3M dataset for caption contrastive fine-tuning. Additionally, we explore augmenting the training data with the Wikitext-103 dataset (Merity et al. 2016) for MNTP training and the E5 dataset (Springer et al. 2024) as supplementary pure-text data for caption contrastive fine-tuning.
Stage1 Stage2 Avg
I2T T2I
78.3 75.5
Linear($\times 1$) 79.2 76.7
Linear($\times 2$) 80.1 76.8
Linear($\times 4$) 80.4 77.9
Linear($\times 4$) Linear($\times 4$) 80.5 77.7
Transformer($\times 1$) 80.2 77.3
Transformer($\times 1$) Transformer($\times 1$) 80.5 77.3
Table 8: Ablation experiments for adaptor design on Stage 1 caption contrastive finetune and Stage 2 LLM2CLIP post training. The 4-layer Linear Adaptor has 67.1M parameters, while the single layer Transformer Adaptor has 67.6M.
Architecture Design Ablation. We examine various adaptor architecture in Table A7. Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP fine-tuning. Therefore, we default to using no adaptor in Stage 1.
Methods Training Loss Testing Text Encoder Average
I2T T2I
CLIP CL(CLIP-T, CLIP-V) CLIP-T 74.4 72.0
Directly Finetune CL(CLIP-T, CLIP-V) CLIP-T 74.5 72.3
+LLM2CLIP a) CL(LLM, CLIP) LLM 83.9 82.1
b) CL(LLM, CLIP-V)+ CL(CLIP-T, CLIP-V) CLIP-T 74.4 72.0
LLM 83.6 81.8
c) CL(LLM, CLIP-V) +CL(CLIP-T, CLIP-V) +CL(CLIP-T, LLM) CLIP-T 74.0 72.1
LLM 83.7 81.4
d) CL(LLM, CLIP-V) +CL(CLIP-T, CLIP-V)+ CL(Cat(CLIP-T, LLM), CLIP-V) CLIP-T 74.8 71.3
LLM 83.8 82.3
Cat(CLIP-T,
LLM)
84.7 82.8
Table 9: LLM2CLIP Training Method Analysis. We experimented with various possibilities for fine-tuning between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T). ”CL” denotes the contrastive learning loss, and ”Cat” means concatenation. Items a–d correspond to the a–d in Section 3.2.
Training Method Ablation. We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in  Table A5, observing the following key insights: 1. Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential. 2. SimCSE is identified as the most critical loss function. MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance. 3. Supervised SimCSE substantially outperforms unsupervised SimCSE. 4. The performance difference between causal and bidirectional attention is negligible.
Ablation across different LLM backbones. Table A6 compares a variety of text-embedding models plugged into our pipeline. Our approach delivers substantial gains over all contenders. The line highlighted in red shows that plain Llama 3.1-8B—without our caption-contrastive (CC) fine-tuning—performs very poorly and even harms the original CLIP, underscoring the necessity of the CC stage. We also evaluated several state-of-the-art embedding models, including bge, jina-v3, and the VLM2Vec embedding obtained from a multimodal LLM fine-tune. In every case, the LLM2CLIP variant still achieves the best results, confirming the effectiveness of our training recipe.
4.4 Ablation Study: Stage-2 LLM2CLIP
Adaptor Design. As shown in Table A7, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training. Comparing Linear ($\times 4$) and Transformer ($\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure. Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance. Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
Training-method ablation. We explored multiple ways to integrate an LLM during CLIP fine-tuning (implementation details appear in Table A8. Comparing experiments (a)–(b) and (b)–(c) shows that neither (i) optimising two separate losses nor (ii) explicitly aligning the two text encoders brings any meaningful benefit. This observation motivates our choice to replace, rather than reuse, CLIP’s original text encoder in the main pipeline. Experiment (d)—which performs contrastive learning on the concatenated outputs of both encoders—does deliver a noticeable uplift; for the sake of simplicity, we do not adopt it as our default, but regard it as a promising future work.
5 Conclusion
We have revisited the problem of upgrading CLIP with large-language-model knowledge and proposed a lightweight pipeline that first turns an LLM into a discriminative caption-embedding module and then inserts it into CLIP via a cost-effective fine-tuning stage. With only million-scale training pairs and a compute budget essentially identical to standard CLIP fine-tuning, our method delivers sizeable gains—on top of already SOTA baselines—in image–text retrieval, image classification (zero-shot & linear probe), zero-shot/supervised detection/segmentation, and even as the visual encoder inside a multimodal LLM.
Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (NSFC Granted No. 62276190).
Appendix A Reproducibility Statement
To ensure the reproducibility of our work, we are committed to making all the training and testing code, along with the datasets used in our experiments, publicly available in our code repository. Additionally, we will open-source our LLM2CLIP-enhanced versions of SigLip2, EVA02, and OpenAI’s CLIP to facilitate applications within the open-source community requiring high-quality embedding models.
We will provide comprehensive documentation alongside the codebase to facilitate straightforward replication of our experiments. All hyperparameter settings, training configurations, and preprocessing steps will be thoroughly documented and made accessible to the research community.
Additionally, we will include detailed environment setup instructions, dependency requirements, and step-by-step execution guidelines to minimize barriers for reproduction. Our commitment to open science ensures that researchers can easily validate, extend, and build upon our contributions.
Appendix B Models and Checkpoints
Table A2 lists all the models and checkpoints used in our experiments, ensuring transparency and facilitating reproducibility. It includes both vision and text encoders, with corresponding links to their publicly available implementations or weights.
Appendix C Experiment Details
C.1 Training Setup
Settings for Stage-1
We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention. No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding. MNTP is not used by default. Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings (Springer et al. 2024) into our contrastive training process.Our models are trained with LoRA ($r=16$, $\alpha=32$), bfloat16 quantization, gradient checkpointing, and FlashAttention-2 for memory efficiency. Training uses the AdamW optimizer (learning rate $2\times 10^{-4}$, 300-step linear warm-up), sequence length 512, effective batch size 2048, and 1 epoch over 30M(1.5m paired pure-text and captions dataset) samples on 32 NVIDIA A100 GPUs.
Settings for Stage-2
We employed the AdamW optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05. For every training step we randomly sampled a single caption from the DreamLIP caption data. All models were trained for 4 epochs by default. The embedding dimension for both modalities was set to 1280.
C.2 Evaluation
Multilingual Retrieval
We evaluate our models on three cross-modal retrieval benchmarks that support multilingual and cross-lingual evaluation: Flickr-CN, COCO-CN, and XM3600. The results are shown in Table 2 of the main text..
Flickr-CN (Lan et al. 2017) and COCO-CN (Li et al. 2019) are Chinese adaptations of the popular Flickr30K and MS-COCO datasets, respectively. These datasets contain images paired with Chinese captions, enabling evaluation of cross-modal retrieval performance in Chinese.
Crossmodal-3600(XM-3600)
(Thapliyal et al. 2022) is a geographically-diverse multilingual dataset comprising 3,600 images annotated with human-generated reference captions in 36 languages . The images were carefully selected from across the world, covering regions where the 36 languages are natively spoken.
Open-Vocabulary Semantic Segmentation
CLIP-like models enable patch-level semantic matching by aligning visual and textual representations, facilitating open-vocabulary semantic segmentation. Following MaskCLIP (Zhou et al. 2022), we compute cosine similarity between patch embeddings and sentence embeddings to generate segmentation masks.
We evaluated our models on four benchmark datasets: ADE20K (Zhou et al. 2017), COCO-Stuff164k (Caesar et al. 2018), VOC20 (Everingham et al. 2015), and Cityscapes (Cordts et al. 2016) using mean Intersection over Union (mIoU) to measure segmentation accuracy. The results are shown in Table 4 of the main text.
Open-Vocabulary Object Detection
To assess the fine-grained localization capabilities of our approach, we integrate LLM2CLIP as the backbone network for open-vocabulary detection tasks. Following established protocols from prior work (Wu et al. 2023), we adopt a two-stage detection framework for evaluation. We conduct experiments on the OV-COCO benchmark and report box Average Precision (AP) at IoU threshold 0.5 across three categories: base classes (seen during training), novel classes (unseen during training), and all classes combined. The novel class performance particularly demonstrates the model’s generalization ability to previously unseen object categories. The strong results on novel classes, as shown in Table 4 of the main paper, demonstrate our method’s effectiveness in recognizing and localizing previously unseen objects, highlighting its robust generalization capabilities for open-vocabulary detection tasks.
Supervised Instance Segmentation.
To evaluate the dense prediction capabilities of our method on supervised instance segmentation tasks, we employ LLM2CLIP as the backbone for ViTDet (Li et al. 2022b) within the Mask R-CNN (He et al. 2017) framework. The models are trained on the COCO dataset using 4 A100 machines with 2 images per GPU, resulting in a total batch size of 64. We train for 25 epochs using the AdamW optimizer with a learning rate of 0.0001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.1.
Our approach achieves results with 45.6 mAP for object detection and 38.7 mAP for instance segmentation, as shown in Table 4 of the main text, demonstrating the effectiveness of our cross-modal pre-training in improving dense prediction tasks. These results highlight the model’s ability to transfer learned representations to complex downstream tasks requiring both object localization and precise pixel-level segmentation.
Comparisons on General Multimodal Benchmarks.
We evaluate LLM2CLIP (EVA02-L-336) as a visual feature extractor for multimodal large language models, comparing it against the standard CLIP baseline. Our experiments are conducted using LLaVA-v1.5-7B, which originally employs CLIP as its visual encoder. To ensure fair and rigorous comparison, we maintain all parameter configurations identical to the original LLaVA setup and utilize the same training data provided by the LLaVA framework. The comprehensive evaluation results on various multimodal understanding benchmarks((i.e., VQA (Goyal et al. 2017) (Hudson and Manning 2019) (Gurari et al. 2018) (Lu et al. 2022) (Singh et al. 2019), POPE (Li et al. 2023b), MMBench (Liu et al. 2024), SEED (Li et al. 2023a)) are presented in Table 5 of the main text.
Figure A1: Comprehensive evaluation results on cross-modal retrieval tasks. The radar plot illustrates the performance of different models across multiple benchmarks. Our methods consistently achieve higher scores across all evaluation metrics compared to baseline models.
Appendix D More Analysis
Quantitative Analysis of the Original LLM’s Feature Separability.
As shown in Table A1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs). We observed that large language models, such as Llama 3 8B, achieve very low accuracy, struggling to distinguish effectively between different captions of the same image. This result highlights the necessity of the embedding-oriented training (caption contrastive fine-tuning) explored in this work for adapting LLMs specifically for the CLIP scenario.
Language Model Top 1 Acc.
CLIP-L/14 25.2
EVA02-L/14 27.1
Llama3-8B 5.2
Llama3.2-1B 5.6
Llama3-8B-$C$ 29.5
Llama3.2-1B-$C$ 29.4
Table A1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO (Lin et al. 2014) 5K test set.
Model Link
Vision Encoders
ViT-B/16 https://huggingface.co/openai/clip-vit-base-patch16
ViT-L/14 https://huggingface.co/openai/clip-vit-large-patch14
ViT-L/14-336 https://huggingface.co/openai/clip-vit-large-patch14-336
EVA-L-224 https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02˙CLIP˙L˙psz14˙s4B.pt
EVA-L-336 https://huggingface.co/QuanSun/EVA-CLIP/blob/main/EVA02˙CLIP˙L˙336˙psz14˙s6B.pt
SigLIP SO/14 https://huggingface.co/google/siglip-so400m-patch14-224
SigLIP2 SO/14 https://huggingface.co/google/siglip2-so400m-patch14-224
Text Encoders
bge-en-icl https://huggingface.co/BAAI/bge-en-icl
LLM2Vec-Llama-3-8B https://huggingface.co/McGill-NLP/LLM2Vec-Meta-Llama-3-8B-Instruct-mntp
NV-Embed-v2 https://huggingface.co/nvidia/NV-Embed-v2
VLM2Vec https://huggingface.co/TIGER-Lab/VLM2Vec-LLaVa-Next
bge-m3-XLM-R https://huggingface.co/BAAI/bge-m3
jina-v3-XLM-R https://huggingface.co/jinaai/jina-embeddings-v3
e5 https://huggingface.co/intfloat/multilingual-e5-large
Qwen2.5-0.5B https://huggingface.co/Qwen/Qwen2.5-0.5B
Llama-3.2-1B https://huggingface.co/meta-llama/Llama-3.2-1B
Llama-3-8B https://huggingface.co/meta-llama/Meta-Llama-3-8B
DeepSeek-R1-Distill-Llama-8B https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B
Llama-3.1-8B https://huggingface.co/meta-llama/Llama-3.1-8B
Table A2: Models and Checkpoints used in our experiments, including both vision and text encoders.
(a) Multilingual zero-shot image-to-text retrieval (Recall@1)
(b) Multilingual zero-shot text-to-image retrieval (Recall@1)
Figure A2: Multilingual zero-shot retrieval results on Crossmodal-3600.
@@ Appendix D More Analysis
Table A3: Ablation study on the proportion of dense captions from different MLLMs. Here, we explore the impact of both original short image captions and synthetic dense captions generated by MLLMs on the performance of LLM2CLIP.
@@
Llama-3.1-8B-$C$ LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
Table A4: Efficiency study of LLM2CLIP. We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size. We report training hours, batch size, and retrieval accuracy for each experiment. During offloading, we precompute text captions offline into embeddings using the LLM. This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
Stage1 Flickr COCO ShareGPT4V Urban-1k DOCCI Avg
@@
Table A5: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.
Method Flickr COCO ShareGPT4V Urban-1k DOCCI Avg
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
bge-m3-XLM-R 80.7 70.3 51.4 42.0 84.5 86.5 56.8 63.4 51.7 55.9 65.0 63.6
jina-v3-XLM-R 84.4 73.7 56.3 45.6 90.0 90.9 70.8 74.1 66.7 70.6 73.6 71.0
e5 (XLM-R) 86.4 75.3 56.4 46.2 88.4 88.5 71.1 77.3 67.5 71.2 74.0 71.7
VLM2VEC 91.6 79.8 61.3 51.5 93.9 91.0 90.9 92.0 80.5 86.0 83.6 80.1
VLM2VEC (finetune) 90.2 79.3 60.0 50.1 89.8 91.4 76.9 85.8 74.1 78.7 78.2 77.1
Qwen2.5-0.5B-$C$ 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
Llama-3.2-1B-$C$ 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Llama-3-8B-$C$ 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
DeepSeek-R1-Distill-Llama-8B-$C$ 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
Llama-3.1-8B-$C$ 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
Table A6: Ablation for using different text encoder. BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs. The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
Linear($\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
Linear($\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
Linear($\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Linear($\times 4$) Linear($\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
Transformer($\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
Transformer($\times 1$) Transformer($\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
Table A7: Ablation experiments for adaptor design on Stage 1 caption contrastive finetune and Stage 2 LLM2CLIP post training. In parentheses, ($\times\textit{l}$) denotes an adaptor with $\times\textit{l}$ layers. The 4-layer Linear Adaptor has 67.1M parameters, while the single layer Transformer Adaptor has 67.6M parameters. This experiment is conducted using llama 3.1 1B and OpenAI CLIP ViT-L 224, with CC 3M data utilized in both Stage 1 and Stage 2. The light blue row indicate our default setting.